Required information Problem 6-43 (LO 6-2) [The following information applies to the questions displayed below.] Calvin reviea. Calvins AG is $290,000. Answer is complete and correct. Charitable contribution deduce Carryover $ 110,140b. Calvins AGI is $290,000 but the State Museum told Calvin that it plans to sell the painting. Charitable contribution dedu

c. Calvins AGI is $144,000. Charitable contribution deduction Carryoverd. Calvins AG is $290,000 and Hobbs is a private nonoperating foundation. Charitable contribution deduction Carryovere. Calvins AGI is $290,000 but the painting is worth $124,000. Charitable contribution deduction Carryover
Answer #1

a. Calvin can deduct $ 110,140. All the contributions are deductible except the donation to the needy family.This donation will not qualify for a charitable deduction because the family is not a qualified charity.In contrast, a donation of food and clothes to a qualifying organization, such as the Red Cross, would qualify. The IBM stock is capital gain property, so Calvin can deduct the FMV of the stock i.e. $ 60,000 subject to 30 percent of AGI i.e. $ 290,000 * 30% = $ 87,000 ceiling. The painting is not capital gain property because it has not appreciated in value. Hence, Calvin can only deduct the value of the painting subject to the 50 percent of AGI ceiling i.e. $ 290,000 * 50% = $ 145,000. Calvin’s cumulative donations are $110,140, which does not approach the 50 percent limit.

Particular Amount
Charitable Contribution Deduction $ 110,140
Carry Over $ 0

b. No difference with a) answer because the painting is not capital gain property.

Particular Amount
Charitable Contribution Deduction $ 110,140
Carry Over $ 0

c.The IBM stock is capital gain property and the painting is not capital gain property. Hence, Calvin can only deduct the value of the stock subject to the lesser of :

(1) the value of the stock up to the 30 percent AGI limit i.e. $ 144,000 *30% = $ 43,200, or

(2) the remaining amount of deduction to reach the 50 percent limit i.e. [($ 144,000 * 50%) - $ 4,140 - $ 46,000] = $ 21,860

Hence, Calvin can deduct $72,000 this year consisting of cash of $46,000, painting of $4,140, and the stock of $ 21,860. The remaining value of the stock $ 38,140 ($60,000 - $ 21,860) is carried over to next year subject to AGI limit.

Particular Amount
Charitable Contribution Deduction $ 72,000
Carry Over $ 38,140

d. The IBM stock is capital gain property but because the donee is a private non operating foundation, the deduction for the value of the stock is subject to a 20 percent of AGI limitation.

Hence, Calvin can deduct the lesser of

(1) the value of the stock up to the 20 percent AGI limit i.e $ 290,000 * 20% = $ 58,000 or

(2) the remaining amount of deduction to reach the 50 percent limit i.e ($ 290,000 * 50% -$ 4140 - $ 46,000) =$ 94,860. Hence, Calvin can deduct $ 108,140 this year consisting of cash of $46,000, painting of $4,140, and the stock of $58,000.The remaining value of the stock $2,000 ($60,000 -$58,000) is carried over to next year subject to the 20 percent of AGI limit.

Particular Amount
Charitable Contribution Deduction $ 108,140
Carry Over $ 2,000